At work one day, someone received eleven packages. Speedy delivered four times as many as Ralph's express, while Ralph's express delivered one more than Speedquick service. How many packets did each service deliver to the person?
Answer by lauramg012(4)

Speedy=x
Ralph's express=y
Speedquick service=z
x+y+z=11
x=4y
y=z+1
z=y-1
substitute: 4y+z+1+y-1=11
combine like
terms: 5y+z=11
substitute: 5y+y-1=11
combine like
terms: 6y-1=11
add: 6y=12
divide: y=2
Now you know that Ralph's express delivered 2 packages. If Speedy delivered four times as many packages as Ralph's express, then they delivered 2x4=8 packages. Speedquick service delivered one less than Ralph's express, so they delivered 2-1=1 packages.
Speedy delivered 8 packages; Ralph's express delivered 2 packages; Speedquick delivered 1 package for a total of 11 packages.